# A Matrix Model Proposal for Quantum Gravity and the Quantum Mechanics of Black Holes

### v2

@@ 3 Quantum black hole as fuzzy sphere with half-filled Fermi sea
where we have labeled the oscillators with $a:=(p,k)$. Physically, we may give an interpretation to (3.18) by saying that the fuzzy sphere has been divided into $N^{2}$ cells with unit cell area of $\Delta A=4\pi l_{P}^{2}$ of the Planck size. Each cell, whose location is labeled by $a=(p,k)$, is populated by a pair of oscillators $\xi_{a},\chi_{a}$ which describes the quantum fluctuations over the fuzzy sphere. This is precisely the description of “partons” suggested in Susskind’s holographic picture [8] of the black hole. It is quite amusing that this conjectured feature of holography get a concrete realization in our quantum mechanics of spacetime.
@@
Let us consider the $n=0$ energy state which corresponds to having half of the oscillator levels in the Fermi sea filled. In this case, the total energy (3.18) of the fuzzy sphere system is
$E=\frac{NM_{P}}{2}.$ (3.19)
Using the relations (3.6) and (3.17), this becomes
$E=\frac{R}{2G}.$ (3.20)
@@
This is precisely the Schwarzschild mass-radius relation if the fuzzy sphere is identified with the black hole horizon and the total energy of the system is identified with the mass $M$ of the black hole
@@
Next let us consider the microstates counting. In fact, the level $n$ eigenvalue has a degeneracy of
@@
$S=2N^{2}=\frac{A}{4G},$ (3.24)
which is precisely the Bekenstein-Hawking entropy of a Schwarzschild black hole. It is remarkable that our quantum mechanical theory (2.4) admits solution that matches precisely the desired properties of a quantum black hole. We thus propose that a quantum black hole is described in our theory by a fuzzy sphere geometry in the energy state $n=0$. We note that the state $n=0$ is singled out by the fact that $\Omega_{n}$ is maximized at $n=0$. It is amazing that a quantum black hole in our description is characterized to be one, given the mass of the black hole is fixed, having the maximal allowed amount of microstates. In the next section, we provide further justification by showing that the fuzzy spheres interact with each other in consistent with Newton gravity in the large distance.
@@
4 Multiple black holes and Newton gravity
The reproduction of the static properties of a quantum black hole is a strong evidence that the fuzzy sphere solution does describe a quantum black hole and that our theory (2.4) is a creditable proposal. Below we provide further justification to our proposal by showing that the fuzzy spheres interact in manner that is characteristic of Newton gravity.
@@ 4 Multiple black holes and Newton gravity
$H_{\rm i}^{0}:=\alpha_{\rm i}^{2}\mbox{tr}_{\rm i}\left(-\frac{M_{P}}{N_{\rm i}^{2}}[X_{\rm i}^{0a},X_{\rm i}^{0b}]^{2}-\frac{4M_{P}}{N_{\rm i}^{2}}X_{\rm i}^{0a2}+\frac{M_{P}}{N_{\rm i}^{2}}\mbox{tr}_{\rm i}\psi^{\dagger}\sigma^{a}X_{\rm i}^{0a}\psi\right)+N_{\rm i}M_{P}$ (4.8)
@@
$H^{F}_{\rm i}:=\frac{M_{P}}{N^{2}}\sum_{m,n=1}^{N_{\rm i}}\psi^{\dagger}_{mn}\sigma^{a}x_{\rm i}^{a}\psi_{mn}$ (4.9)
@@
are obtained from the abelian $U(1)$ part $x^{a}_{\rm i}$ of the configuration. In the above, the trace $\mbox{tr}_{\rm i}$ is taken over the $N_{\rm i}\times N_{\rm i}$ subspace. Finally the terms $H_{\rm i}^{F},H_{12}$ represent interaction between the two blocks and depend on their separation distance (in dimensionless matrix units)
@@
This comes because (4.3) implies that $x_{1}^{a}=\alpha_{2}(x_{1}^{a}-x_{2}^{a})$ etc. As a result, the fermionic interaction term $H^{F}_{\rm i}$ is proportional to $\Delta x$, and the bosonic interaction term is proportional to $\Delta x^{2}$
@@
where $\xi^{p}_{k},\chi^{p}_{l}$ are defined as in (3.11) using the eigendecomposition (3.10) for $\sigma^{a}X_{\rm i}^{0a}$ and the indices $m,n$ range over the block of $X_{\rm i}$. Consider the set of states where half of the oscillators over the fuzzy sphere are excited (i.e. $r+s=N_{\rm i}^{2}$), the fuzzy sphere has the energy
$H^{0}_{\rm i}=\frac{M_{P}N_{\rm i}}{2}(2-\alpha_{\rm i}^{2}).$ (4.15)
@@
Next let us consider the interaction term $H^{a}_{\rm i}$. For small $\Delta x/N$, we can consider $H^{a}_{\rm i}$ to be a perturbation and use the perturbation theory for degeneracy eigenstates to determine its correction to the energy. This requires the knowledge of the matrix elements $\langle i|H^{F}_{\rm i}|j\rangle$ and it’s eigenvalues $\lambda_{p}$. Here $|i\rangle,i=1,\cdots,\Omega_{0}:=2^{2N_{\rm i}^{2}}$ are the unperturbed states (4.14) with $r+s=N_{\rm i}^{2}$. This is however not only impossible to do as it involves too large a number of states, but it is also unnecessary. A more meaningful quantity is the ensemble average of the corrections,
$\frac{1}{\Omega_{0}}\sum_{p}\lambda_{p}$ (4.16)
We show in the appendix that this vanishes. As a result, the total energy of the two fuzzy spheres system is given by
$E=\frac{NM_{P}}{2}(1+3\alpha_{1}\alpha_{2})-\frac{4M_{P}}{N}\alpha_{1}\alpha_{2}\Delta x^{2}.$ (4.17)
Here we have used $\alpha_{1}^{3}+\alpha_{2}^{3}=1-3\alpha_{1}\alpha_{2}$. The second term is a correction to the first term in the leading ordering of small $\Delta x^{2}/N^{2}$. In principle, one should include the higher order perturbation as well as the interaction term from the off diagonal blocks of (4.1). By summing over all these contributions, one can expect the final result for the system energy to be a function of $\Delta x/N$ whose leading term in the small $\Delta x/N$ expansion gives (4.17). In other words, the large-$N$ summation would result in the energy for the two fuzzy spheres system
$E=\frac{NM_{P}}{2}(1+5\alpha_{1}\alpha_{2})-NM_{P}\alpha_{1}\alpha_{2}f(\frac{\Delta x}{N}),$ (4.18)
where $f$ is some function with the property that $f(x)=1+4x^{2}+O(x^{3})$ for small $x$. Defining the coordinate distance between the two black holes as $r=\Delta x\,l_{P}$, we have $H=M_{1}+M_{2}+5\mu-\frac{GM_{1}M_{2}}{r}g(r),$ where $M_{\rm i}=N_{\rm i}M_{P}/2$ is the mass of the black hole in isolation, $\mu=M_{1}M_{2}/(M_{1}+M_{2})$ is the reduced mass, and $g(r):=\frac{4r}{R}f(\frac{r}{R})$. Note that the factor $GM_{1}M_{2}$ appears exactly the way one expects for Newton gravity. One can define the interaction energy as $V(r):=H(r)-H(r=\infty)=-\frac{GM_{1}M_{2}}{r}g(r)$. Note that it is generally true that
$V(0)=-\frac{4GM_{1}M_{2}}{R}.$ (4.19)
This is independent of the form of $f$. The removal of the infinite interaction energy at the origin is generic and is a prediction of our proposed theory of quantum spacetime. Note also that the Newton gravity
@@
will be recovered in the leading order of large distance $r/R=\Delta x/N\gg 1$ if $f$ is such that $g=1+O(1/r)$. An example is $f(x)=(1-x)/(1-x-4x^{2})$. In this case, $g(r)=(r/R-1)/(r/R+\frac{1}{4}(R/r-1))$. It is important to devise method to reliably compute the potential $V(r)$ between the fuzzy spheres in order to confirm that gravity does emerge from our proposed theory of quantized spacetime. More generally it is important to understand how a description involving metric and Einstein gravity emerges in the low energy approximation.

### v3

@@ 3 Fuzzy sphere solution with Fermi sea
Physically, (3.17) may be given an interpretation that the fuzzy sphere has been divided into $N^{2}$ cells with unit cell area of $\Delta A=4\pi l_{P}^{2}$ of the Planck size. Each cell, whose location is labeled by $a=(p,k)$, is populated by a pair of oscillators $\xi_{a},\chi_{a}$ which describes the quantum fluctuations over the fuzzy sphere. This is strikingly similar to the description of “partons” in the holographic picture of black hole suggested by Susskind [8]. It is quite amusing that this conjectured feature of holography get a concrete realization in our quantum mechanics of spacetime. In the next section, we propose to identify the quantum Schwarzschild black hole with the fuzzy sphere solution with a half-filled Fermi sea $n=0$ in our model.
Before we move on, it is necessary to discuss the stability of the fuzzy sphere solution. In general relativity, it is known that the Schwarzschild metric and the Kerr metric is stable against mode perturbations that do not change the mass $M$ and angular momentum $J$ of these vacuum solutions. This statement of mode stability was originally proven [32, 33] for perturbations that can be written as a superposition of spherical harmonic modes. Recently, it has been shown that [34] Schwarzschild metric is stable against general linear perturbations that does not necessarily need to be written as a finite sum of spherical harmonic modes. However, the linear stability of Kerr metric remains an open problem. The full non-linear stability of vacuum solution in general relativity is a hard problem and has only been established for the Minkowski metric [35]. For us here, we will be satisfied with a mode analysis and shows that our fuzzy sphere solution is stable against mode perturbations that do not change the energy of the solution.
For linearized analysis, we look at the quadratic potential of the fluctuations $\delta X^{a}$. Over a general classical solution $X^{a}$, we have
$U=\frac{M_{P}}{N^{2}}V,\quad V=-\mbox{tr}\left(F_{ab}^{2}+2[X^{a},X^{b}][\delta X^{a},\delta X^{b}]+4\delta X^{a}\delta X^{a}\right),$ (3.18)
where
$F_{ab}:=[X^{a},\delta X^{b}]-[X^{b},\delta X^{a}]$ (3.19)
is the field strength. The diagonalization of $V$ is standard. It is convenient to introduce the derivative operator $L^{a}$ whose action on a matrix $f$ is given by $L^{a}f=[X^{a},f]$. In terms of which we have
$F_{ab}=-i\epsilon_{abc}(\varepsilon\cdot L)_{cd}\delta X^{d},$ (3.20)
where $\varepsilon_{a}$ is a vector matrices with components $(\varepsilon_{a})_{bc}:=-i\epsilon_{abc}$. So far this is general. For $X^{a}$ given by the fuzzy sphere (3.3), (3.4), the second term in $V$ of (3.18) can also be expressed in terms of $F_{ab}$ and the potential takes the compact form
$V=2\mbox{tr}(\delta X^{a}N_{ab}\delta X^{b}),\quad\mbox{where}\quad N_{ab}:=\left((\varepsilon\cdot L)^{2}-(\varepsilon\cdot L)-2\right)_{ab}.$ (3.21)
As a result, the equation of motion for the linearized fluctuation is given by
$\delta\ddot{X}^{a}+\frac{4a_{0}^{2}M_{P}^{2}}{N^{2}}N_{ab}\delta X^{b}=0,$ (3.22)
and the study of the mode stability of the fluctuations reduces to the eigenvalue problem of $N_{ab}$:
$N_{ab}\phi^{b}=\lambda\phi^{a}.$ (3.23)
The operator $N_{ab}$ is diagonalized by the vector harmonics. To construct them, note that $\varepsilon_{a}$ is an angular momentum of angular momentum quantum number $\ell_{\varepsilon}=1$ since
$\varepsilon_{a},\varepsilon_{b}]=i\epsilon_{abc}\varepsilon_{c},\quad\varepsilon_{a}^{2}=2.$ (3.24)
An eigenbasis is given by $|\ell_{\varepsilon},m_{\varepsilon}\rangle$, $m_{\varepsilon}=-1,0,+1$:
$\varepsilon^{2}|\ell_{\varepsilon},m_{\varepsilon}\rangle=2|\ell_{\varepsilon},m_{\varepsilon}\rangle,\quad\varepsilon_{z}|\ell_{\varepsilon},m_{\varepsilon}\rangle=m_{\varepsilon}|\ell_{\varepsilon},m_{\varepsilon}\rangle.$ (3.25)
Note also that for the fuzzy sphere background, $L^{a}$ is an angular momentum operator. An eigenbasis is given by $|\ell,m_{z}\rangle$,
$L^{2}|\ell,m_{z}\rangle=\ell(\ell+1)|\ell,m_{z}\rangle,\quad L_{z}|\ell,m_{z}\rangle=m_{z}|\ell,m_{z}\rangle$ (3.26)
for $m_{z}=-\ell,\cdots,\ell$ and $0\leq\ell\leq N-1$. Here the angular momentum quantum number $\ell$ is cut off by $N$ due to the noncommutativity of the fuzzy sphere. The spherical harmonics ${\hat{Y}}^{\ell}_{m_{z}}$ is now obtained in the matrix representation as $({\hat{Y}}^{\ell}_{m_{z}})_{n_{1}n_{2}}=\langle n_{1}n_{2}|\ell,m_{z}\rangle$. The eigenvalue problem of the operator $\varepsilon\cdot L$ can now be easily solved with the help of the total angular momentum operator $J^{a}:=\varepsilon^{a}+L^{a}$. $(J^{2},J_{z})$ is diagonalized by
$J^{2}|J,M_{J}\rangle=J(J+1)|J,M_{J}\rangle,\quad J_{z}|J,M_{J}\rangle=m_{J}|J,M_{J}\rangle,$ (3.27)
and
[…]
As a result, $\varepsilon\cdot L=\frac{1}{2}(J(J+1)-\ell(\ell+1)-2)$ and $\lambda$ has the eigenvalues,
$\varepsilon\cdot L=\begin{cases}0&\\ -(\ell+1)&\\ -1&\\ \ell&\\ \end{cases},\quad\lambda=\begin{cases}-2&\qquad\mbox{for $\ell=0,J=1$}\\ \ell(\ell+3)&\qquad\mbox{for $\ell\geq 1,J=\ell-1$}\\ 0&\qquad\mbox{for $\ell\geq 1,J=\ell$}\\ (\ell+1)(\ell-2)&\qquad\mbox{for $\ell\geq 1,J=\ell+1$}\\ \end{cases}.$ (3.29)
For given $\ell$ and $J$ as given by (3.28), the corresponding eigenstate can be constructed by using the CG coefficient $C^{JM_{J}}_{\ell m_{z}1p}$
$|J,M_{J}\rangle=\sum_{m_{z}=-\ell}^{\ell}\;\;\sum_{p=0,\pm 1}C^{JM_{J}}_{\ell m_{z}1p}|\ell m_{z}\rangle\otimes|1p\rangle,$ (3.30)
where here we have used $p=0,\pm 1$ to denote the polarization. The eigenfunction $\phi^{a}$ of (3.23) can then be constructed with the matrix elements,
$(\phi^{a})_{n_{1}n_{2}}=(\langle n_{1}n_{2}|\otimes\langle a|)|JM_{J}\rangle,$ (3.31)
where $|n_{1}n_{2}\rangle$ are the basis states of the matrix space and $|a\rangle$ are the basis states of the 3-dimensional vector space. As a result, we obtain the eigenvector $\phi^{a}$ as given by the vector spherical harmonics
$\phi^{a}=\sum_{m_{z}=-\ell}^{\ell}\;\;\sum_{p=0,\pm 1}C^{JM_{J}}_{\ell m_{z}1p}\;e^{p}_{a}\;{\hat{Y}}^{\ell}_{m_{z}}:=({\hat{Y}}^{JM_{J}}_{\ell})_{a},$ (3.32)
where $e^{p}_{a}:=\langle a|1p\rangle$ is the $a$-th component of the polarization vector $e^{p}$. Note that the eigenvalue (3.29) is independent of $M_{J}$ and so each of the eigenvalues in (3.29) has a degeneracy of $2J+1$. In total, we have $3N^{2}$ eigenmodes in (3.29). Not all of these modes are admissible fluctuations, however.
The perturbation is unstable if the eigenvalue is negative, i.e. for the modes $\ell=0,J=1$ and $\ell=1,J=2$. The mode $\ell=1,J=2$ corresponds to a scaling $\delta X^{a}\propto X^{a}$ of the fuzzy sphere. This changes the energy of the fuzzy sphere. The mode $\ell=0,J=1$ corresponds to a shift of an overall $U(1)$ factor, which is not allowed since $X^{a}$ is traceless. Fluctuations represented by the other modes, energy preserving or not, are stable. Therefore we find that the single fuzzy sphere solution is stable against linearized energy preserving perturbations.
Finally, we note that upon quantization, the quadratic fluctuation modes become oscillators with frequency determined by $\lambda$ in (3.29). We adopt a definition of the quantized bosonic Hamiltonian by subtracting a constant such that the fuzzy sphere has vanishing zero point energy. In other word, the ground state of the single fuzzy sphere has a vanishing zero point energy for both the bosonic as well as the fermionic oscillations.
4 Quantum black hole as fuzzy sphere with a half-filled Fermi sea
@@
Consider the fuzzy sphere solution with a half-filled Fermi sea ($n=0$). The total energy of the fuzzy sphere system is given by
$E=\frac{\gamma NM_{P}}{2}.\quad\mbox{where $\gamma:=a_{2}+2a_{3}-1$}.$ (4.1)
To compare with the energy of the Schwarzschild black hole, let us introduce dimensional coordinates $Y^{a}=2l_{P}X^{a}$ for some length scale $l_{P}$, the fuzzy sphere solution becomes
$Y^{a},Y^{b}]=\frac{2iR}{\sqrt{N^{2}-1}}\epsilon_{abc}Y^{c},\quad\sum_{a}Y^{a2}=R^{2}{\bf 1},$ (4.2)
where $R^{2}=(N^{2}-1)l_{P}^{2}$. This describes a fuzzy sphere of radius $R=Nl_{P}$ in the large $N$ limit. The energy (4.1) can then be put in the form
$E=\frac{R}{2G}$ (4.3)
if we identify $M_{P}$ and $l_{P}$ with the Newton constant $G$ as
$G=\frac{l_{P}}{\gamma M_{P}}.$ (4.4)
@@
The relation (4.3) is precisely the Schwarzschild mass-radius relation if the fuzzy sphere is identified with the black hole horizon and the total energy of the system is identified with the mass $M$ of the black hole
@@
Next let us consider the microstates counting. The level $n$ eigenvalue has a degeneracy of
@@
$S=2N^{2}.$ (4.8)
This is precisely the Bekenstein-Hawking entropy of a Schwarzschild black hole if
$l_{P}=\sqrt{\frac{2G}{\pi}}.$ (4.9)
As a result of (4.4), $M_{P}$ is given by
$M_{P}=\frac{1}{\gamma}\sqrt{\frac{2}{\pi G}}.$ (4.10)
It is remarkable that our quantum mechanical theory (2.2) admits solution that matches precisely the desired properties of a quantum black hole. We thus propose that a quantum black hole is described in our theory by a fuzzy sphere geometry with a half-filled fermi sea. We note that the matching works for any coefficients $a_{0},a_{2},a_{3}$ as long as $\gamma>0$. This arbitrariness of the action can be fixed by considering the rotating Kerr black hole [36]. We note that the state $n=0$ is singled out by the fact that $\Omega_{n}$ is maximized at $n=0$. It is amazing that a quantum black hole in our description is characterized to be one, given the mass of the black hole is fixed, having the maximal allowed amount of microstates. In the next section, we provide further justification by showing that the fuzzy spheres interact with each other with a $GM_{1}M_{2}$ dependence that is characteristic of Newton gravity.
@@
5 Multiple black holes
The reproduction of the static properties of a quantum black hole supports the conjecture that the fuzzy sphere solution does describe a quantum black hole and that the theory (2.2) is a creditable proposal. Below we perform a preliminary analysis on the interaction of fuzzy spheres and show that some characteristic features of Newton gravity can be reproduced.
@@ 5 Multiple black holes
$H_{\rm i}^{0}:=\alpha_{\rm i}^{2}\mbox{tr}_{\rm i}\left[-\left(\frac{M_{P}}{N_{\rm i}^{2}}[X_{\rm i}^{0a},X_{\rm i}^{0b}]^{2}+\frac{4M_{P}}{N_{\rm i}^{2}}X_{\rm i}^{0a2}\right)+\frac{a_{2}M_{P}}{N_{\rm i}^{2}}\mathop{:}\psi^{\dagger}\sigma^{a}X_{\rm i}^{0a}\psi\mathop{:}\right]+a_{3}r_{X_{\rm i}}M_{P}$ (5.8)
@@
$H^{\prime}_{\rm i}=\frac{a_{2}M_{P}}{N^{2}}\sum_{m,n=1}^{N_{\rm i}}\mathop{:}\psi^{\dagger}_{mn}\sigma^{a}x_{\rm i}^{a}\psi_{mn}\mathop{:}$ (5.9)
@@
are obtained from the abelian $U(1)$ part $x^{a}_{\rm i}$ of the configuration. In the above, the trace $\mbox{tr}_{\rm i}$ is taken over the $N_{\rm i}\times N_{\rm i}$ subspace. Finally the terms $H_{\rm i}^{\prime},H_{12}$ represent interaction between the two blocks and depend on their separation distance (in dimensionless matrix units)
@@
This comes because (5.3) implies that $x_{1}^{a}=\alpha_{2}(x_{1}^{a}-x_{2}^{a})$ etc. As a result, the fermionic interaction term $H^{\prime}_{\rm i}$ is proportional to $\Delta x$, and the bosonic interaction term is proportional to $\Delta x^{2}$
@@
where $\xi^{p}_{k},\chi^{p}_{l}$ are defined as in (3.8) using the eigendecomposition (3.7) for $\sigma^{a}X_{\rm i}^{0a}$ and the indices $m,n$ range over the block of $X_{\rm i}$. Consider the set of states where half of the oscillators over the fuzzy sphere are excited (i.e. $r+s=N_{\rm i}^{2}$) the fuzzy sphere has the energy
$H^{0}_{\rm i}=\frac{M_{P}N_{\rm i}}{2}(a_{2}+2\alpha_{\rm i}a_{3}-\alpha_{\rm i}^{2})$ (5.15)
@@
Next, let us consider the interaction term $H^{a}_{\rm i}$. For small $\Delta x/N$, we can consider $H^{a}_{\rm i}$ to be a perturbation and use the perturbation theory for degeneracy eigenstates to determine its correction to the energy. This requires the knowledge of the matrix elements $\langle i|H^{\prime}_{\rm i}|j\rangle$ and it’s eigenvalues $\lambda_{p}$. Here $|i\rangle,i=1,\cdots,\Omega_{0}:=2^{2N_{\rm i}^{2}}$ are the unperturbed states (5.14) with $r+s=N_{\rm i}^{2}$. This is however not only impossible to do as it involves too large a number of states, Physically, since the set of microstates is not observed, it is more meaningful to consider the average corrections over the set of microstates:
$\langle H^{\prime}_{\rm i}\rangle:=\sum_{|i\rangle\in\mathbb{V}_{0}}p_{i}\langle i|H^{\prime}_{\rm i}|i\rangle,$ (5.16)
where $p_{i}$ is the probability of occurrence of the state $|i\rangle$. For an isolated system, we can take the probability to be equal for each of the state and hence $p_{i}=1/\Omega_{0}$. This corresponds to a microcanonical ensemble. We show in the appendix that this ensemble average is given by
$\langle H^{\prime}_{\rm i}\rangle=\frac{a_{2}M_{P}}{N^{2}}\mbox{tr}(\sigma^{a}X_{\rm i}^{0a}\sigma^{b}x_{\rm i}^{b})=0.$ (5.17)
As a result, the total energy of the two fuzzy spheres system is given by
$E=\mbox{const.}-\frac{4a_{2}M_{P}}{N}\alpha_{1}\alpha_{2}\Delta x^{2}.$ (5.18)
where the constant term is independent of $\Delta x^{2}$. The second term is a correction to the first term in the leading ordering of small $\Delta x^{2}/N^{2}$. In general, one should include the higher order perturbation as well as the interaction term from the off diagonal blocks of (5.1) and quantum loop effects [39]. By summing over all these contributions, one can expect the final result for the energy of the two fuzzy spheres system to take the form
$E=\mbox{const.}-\gamma NM_{P}\alpha_{1}\alpha_{2}f(\frac{\Delta x}{N}),$ (5.19)
where $f$ is some function of $\Delta x/N$. It takes the form $f(x)=1+\frac{4a_{2}}{\gamma}x^{2}+O(x^{3})$ for small $x$ in our tree level computation. Defining the coordinate distance between the two black holes as $r=\Delta x\,l_{P}$, and define the interaction energy as $V(r):=E(r)-E(r=\infty)$. We have $V(r)=-\frac{GM_{1}M_{2}}{r}g(r)$ where $M_{\rm i}=\gamma N_{\rm i}M_{P}/2$ is the mass of the black hole in isolation, and $g(r):=\frac{4r}{R}f(\frac{r}{R})$. Note that the factor $GM_{1}M_{2}$ appears exactly the way one expects for Newton gravity. In order to reproduce Newton gravity
@@
in the large distance limit $r/R=\Delta x/N\gg 1$, the function $f$ has to be such that $g=1+O(1/r)$. In order to check whether Newton gravity arises in the large distance limit, it is important to devise method to reliably compute the potential $V(r)$ between the fuzzy spheres. We note that in our model a classical force arises between the fuzzy sphere due to the mass term. It is also possible to consider a variant of our model by considering a Chern-Simons term instead of a mass term, see. e.g. [40, 41, 42] for reduced matrix model with a Chern-Simons term. The model also admits fuzzy spheres as solution but in this case there is no classical force between them and the Newton gravity would have to emerge from the loop, similar to the situation in the BFSS matrix model [21, 39]. We leave this important issue for further analysis. More generally it is important to understand whether and how tensorial Einstein gravity may emerge in the low energy approximation in our model.